OpenStudy (anonymous):

Twelve more than twice a certain number is six fewer than three times the number. What is the number? 6 12 16 18 24

OpenStudy (anonymous):

2x + 12 = 3x - 6 12 + 6 + 2x - 2x = 3x - 2x - 6 + 6 18 = x check: 2(18) + 12 = 3(18) - 6 36 + 12 = 54 - 6 48 = 48 number = 48